Circular No. 9053

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

EDITORIAL NOTICE On MPC 66141, with little explanation, it was announced that the

Minor Planet Center will cease charging for its online publications on 2009 Oct. 1. The “Computer Service” (run since 1983 jointly by the Central Bureau and the MPC) will continue after Oct. 1 as before for paying subscribers, giving access to IAUCs and CBETs.

(136617) 1994 CC M. Brozovic and L. A. M. Benner, Jet Propulsion Laboratory (JPL),

California Institute of Technology (CIT); M. C. Nolan and E. S. Howell, Arecibo Observatory; C. Magri, University of Maine at Farmington; J. D. Giorgini, JPL; P. A. Taylor, Arecibo Observatory; J. L. Margot, University of California, Los Angeles; M. W. Busch, CIT; M. K. Shepard, Bloomsburg University; L. M. Carter, Smithsonian Institution; J. S. Jao, J. Van Brimmer, C. R. Franck, M. A. Silva, M. A. Kodis, D. T. Kelley, and M. A. Slade, JPL; A. Bramson, University of Wisconsin; K. J. Lawrence, JPL; J. T. Pollock, Appalachian State University; P. Pravec, Ondrejov Observatory; and D. E. Reichart, K. M. Ivarsen, J. Haislip, M. C. Nysewander, and A. P. La-Cluyze, University of North Carolina, Chapel Hill, report that Goldstone (8560-MHz, 3.5-cm) and Arecibo (2380-MHz, 12.6-cm) delay-Doppler radar images obtained on June 12–15 reveal that minor planet (136617) is a triple system. Visible range extents in 19-m-resolution Goldstone data from June 12 suggest preliminary diameter estimates of 650 m for the primary and at least 50 m and 100 m for the satellites, whose orbital separations from the primary were at least 0.5 km and 1.2 km. Photometry obtained with the Panchromatic Robotic Optical Monitoring and Polarimetry Telescopes (PROMPT) reveals a 0.05-mag-deep attenuation centered on June 3.04 UT, suggesting a satellite eclipse/occultation, and yields a lightcurve amplitude of 0.09 mag and rotation period of 2.3886 ± 0.0001 hr for the primary.

COMET C/2009 A6 (STEREO) Further to IAUC 9051, another slightly diffuse and somewhat elongated

Kreutz sungrazer has been found on STEREO/SECCHI HI-1A images, reaching peak magnitude ∼ 11.

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 A6 Jan. 2.378 22^{h}11.^{m}1 −14^{o}25′ HI* AW 2009-F31
2009 June 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green